OpenStudy (anonymous):

How do I find the exact value of the expression w/out using a calculator? Sin405 and cos 33 π/4

OpenStudy (anonymous):

sin(405)=sin(360+45)=sin45=1/sqrt(2)

OpenStudy (anonymous):

subtract 360 from 405 and see that it is coterminal with \(405-360=45\) then take \(\sin(45)\)

OpenStudy (anonymous):

cos33 π/4 =cos(32+1)π/4=cos(8π+ π/4)=cosπ/4=1/sqrt(2)

OpenStudy (anonymous):

here we have used sin(360*n+x)=sinx cos(360*n+x)=sinx n being an integer